Imagine you are driving around a horizontal circular track (such as a roundabout at an intersection or freeway exit). You are driving at a constant speed of 9.00 km/h, and your circular path has a radius of 16.0 meters.

a) When you are at the southernmost point on the circular track and driving towards the west, what is your acceleration? Give your answer in m/s2

b) When you are at the easternmost point on the circular track and driving towards the south, what is your acceleration?
Answer #1

The magnitude of Acceleration will be same no matter where you are. Remember, centripetal acceleration always points toward the center of circle.

Here, the acceleration is centripetal acceleration

a = v2 / r

v = 9 km/ h = 2.5 m/s ( 1 km/h = 0.2777 m/s)

Therefore,

a = 2.52 / 16

a = 0.39 m/s2